OpenStudy (anonymous):

An object is dropped at from rest from he top of a 100m building.How long will it take for the object to hit the ground?

OpenStudy (anonymous):

Do you know which equations of motion to use?

OpenStudy (anonymous):

\[\Delta y=Vt+\frac{ 1 }{ 2 } g t ^{2}\]

OpenStudy (anonymous):

Vt = 0 because your initial velocity is 0, g = -9.8 m/s^2, solve for t.

OpenStudy (anonymous):

what does that first part say change in what? what does y stand for?

OpenStudy (anonymous):

y is your change in height. It will be -100m since your initial height is 100m and your final height is 0m.

OpenStudy (anonymous):

i wrote down iv=o t=? d=100m g=9.8m/s^2

OpenStudy (anonymous):

I prefer to use y instead of d in this case because it is a vertical change, not a horizontal.

OpenStudy (anonymous):

So, you have -100m=(0 m/s)(t)-(1/2)(9.8m/s^2)(t^2)

OpenStudy (anonymous):

This simplifies to -100=-4.9t^2

OpenStudy (anonymous):

Divide and then take the square root of 100/4.9. It should be close to the square root of 20.

OpenStudy (anonymous):

i got 2.04 seconds

OpenStudy (anonymous):

That sounds about right. =)

OpenStudy (anonymous):

thanks:)i c in this example it has 4.52 seconds

OpenStudy (anonymous):

And it followed the same "path" we chose, right?

OpenStudy (anonymous):

yea i understand your way but i dnt understand why it has that answer

OpenStudy (anonymous):

What does the question say exactly? Sometimes leaving something small out might change the way the question is answered.

OpenStudy (anonymous):

it has d=vit+1/2at^2 d=1/2 at^2 t=square root 2d/a t=square root 2(100)/9.81=4.52 s

OpenStudy (anonymous):

I just did it out the way I showed you and I got the same answer as the book. I did not get my calculator out before. They way they do it here is instead of multiplying the 9.8 by 0.5 to get 4.9, they multiplied the 100 by 2 instead and then divided by 9.8. The answer works out the same both ways, but I think there may have been an error in the input process on the calculator.

OpenStudy (anonymous):

so you think your answer would still be correct right?

OpenStudy (anonymous):

Compared to what the book gave, I got the same 4.52 seconds.

OpenStudy (anonymous):

What did you put into the calculator?

OpenStudy (anonymous):

D = Vit + 1/2at^2 Starting from rest so Vit can be removed (initial velocity * time) D = 1/2at^2 100m= 1/2at^2 (a = acceleration = gravity = 9.8 m/s^2) 100m=4.9 m/s^2 * t^2 20.408 s^2 = t^2 t = 4.5175s

OpenStudy (anonymous):

Thank you for writing that out greenleaf. =P

OpenStudy (anonymous):

No problem

OpenStudy (anonymous):

k i did 100/24.01=sqrt 4.16......2.04 @greenleaf800073 for you r help and thanks @Hagopian13

OpenStudy (anonymous):

Where did the 24.01 come from?

OpenStudy (anonymous):

Ooooooohhhh, I see.

OpenStudy (anonymous):

You accidentally squared the 4.9 m/s^2. That does not need to squared as the time t is the only part of the equation being squared. That is why your answer worked out to the 2.04 seconds.

OpenStudy (anonymous):

oh ok .u c where it says 100m=4.9m/s^2x t^2....the 20.408 u got from where

OpenStudy (anonymous):

The 20.48 is from 100/4.9.

OpenStudy (anonymous):

yea n the 4.517 is frrom/?

OpenStudy (anonymous):

Taking the square root of 20.48.

OpenStudy (anonymous):

K THANKS ALOTi get it

OpenStudy (anonymous):

Whoop whoop! =)